OpenStudy (anonymous):

Help!!!! I'll give you a Medal Solve, using linear combination. 4x + y = 5 3x + y = 3 A. (0, –3) B. (2, –3) C. (0, 5) D. (–3, 7)

OpenStudy (arabpride):

Put into y = mx + b form 4x + y = 5 -4x -4x >>>y = -4x + 5<<< ~~~~~~~~~~~~~~~~~~ 3x + y = 3 -3x -3x >>>y = -3x + 3<<< ~~~~~~~~~~~~~~~~~~

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

A

OpenStudy (arabpride):

How'd You get your answer @samsteen15

OpenStudy (anonymous):

i did the math

OpenStudy (arabpride):

-.-'' how? you need to provide an explanation...

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

4x+y=5 3x+y=3 plug in A.... 4(0)+(-3)=5 FALSE because, 4*0=0+-3=-3 NOT 5 plug in B 4(2)+(-3)=5 TRUE because, 4*2=8+-3=5 TRUE 3(2)+-3=3 TRUE because 3*2=6+-3=3TRUE so B would be the answer :)

OpenStudy (anonymous):

What one?